OpenStudy (anonymous):

Find the height of a cylinder with a radius of 3.5 cm & surface area of 1000 cm. sq. A)13.42 B)26.83 C)41.97 D)20.0

OpenStudy (lgbasallote):

surface area of a cylinder is \[\LARGE SA = 4\pi r^2 + 2\pi rh\] does that help?

OpenStudy (lgbasallote):

i can explain further if you want....

Parth (parthkohli):

Replace SA with 1000 and r with 3.5

Parth (parthkohli):

Apparently you have to solve for h.

OpenStudy (anonymous):

I DONT GET WHAT EITHER OF YOU ARE SYAING AT ALL & @lgbasallote : YOU HAVE TO FIND THE HEIGHT NOT THE SURFACE AREA

OpenStudy (lgbasallote):

ahh yes...like i said you need only ask..

Parth (parthkohli):

Yeah, we know the surface area and radius, we have to solve for h.

Parth (parthkohli):

\( \color{Black}{\Rightarrow 1000 = 4\pi (3.5)^2 + 2\pi (3.5)h }\) Solve for h, which is the only variable here

OpenStudy (anonymous):

Use the following formula: \[h = (SA - 4\pi r^2)/2\pi r\]

Parth (parthkohli):

@waterineyes Solving for h in the equation is a better start.

OpenStudy (lgbasallote):

now i shall provide you with some magic... \[\large SA = 4\pi r^2 + 2\pi rh\] \[\large SA - 4\pi r^2 = 2\pi r h\] \[\large \frac{SA - 4\pi r^2}{2\pi r} = h\] now can you solve for h?

OpenStudy (lgbasallote):

you will simply substitute SA and r into the equation i provided ;)

Parth (parthkohli):

^ good way to do so

Parth (parthkohli):

\(\Huge \text{<tips cap>}\) yw

OpenStudy (anonymous):

ParthKohli, that is what you think but I don't...

Parth (parthkohli):

@waterineyes unexplained definition isn't the right way, though lgba explained it well enough

OpenStudy (anonymous):

Then writing the direct formula is also not an explanation my dear friend...

OpenStudy (anonymous):

@lgbasallote : I DID THE FORMULA & GOT 4649.5393. WHICH IS NOT CORRECT

Parth (parthkohli):

Thank you for the caps attack, but we don't need one.

OpenStudy (lgbasallote):

did you do \[\huge \frac{1000 - 4(3.14)(3.5)(3.5)}{2(3.14)(3.5)}?\]

Parth (parthkohli):

Calculator FTW!

OpenStudy (anonymous):

yup i did that. and i am using a calculator. why are you being such an a** ? @ParthKohli & maybe i like writing in caps. doesnt seem to bother anyone else !!!

Parth (parthkohli):

Who's being an a**, ma'am?

OpenStudy (anonymous):

you.

Parth (parthkohli):

orly?

OpenStudy (lgbasallote):

try multiplying 4 x 3.14 x 3.5 x 3.5 first what do you get?

OpenStudy (anonymous):

@lgbasallote : 153.86

Parth (parthkohli):

153.86... try dividing that by 2 x 3.14 x 3.5

Parth (parthkohli):

FTW = For the win, not what you think.

OpenStudy (anonymous):

38.49

OpenStudy (anonymous):

@rajathsbhat : not one of my options

OpenStudy (anonymous):

Numerator is : \[1000 - 4(22)(3.5)(3.5)/7\] 1000 - 4(22)(0.5)(3.5) = 1000 - 154 = 846... And the Denominator is :\[2(22)(3.5)/7 = 2(22)(0.5)\] = 22.. Now h = 846/22 = 38.45... I am getting 38.45 as answer guys...

OpenStudy (lgbasallote):

im getting something aroung 38.4-38.5 too

OpenStudy (lgbasallote):

around*

OpenStudy (anonymous):

close to 41.97.

OpenStudy (anonymous):

*close enough*

OpenStudy (lgbasallote):

hah lol..i like your thinking ;)

OpenStudy (anonymous):

It is not close to 41.97.. There is something wrong in the question or options given...

OpenStudy (anonymous):

i have to show my work and your not proving how you get it. . 38.5 is nothing close to 41.97 & @waterineyes : thank you

OpenStudy (anonymous):

Welcome dear...

OpenStudy (lgbasallote):

lol isnt the formula i wrote proof enough why we're getting 38.5 =))

OpenStudy (anonymous):

great point, lgba.

OpenStudy (anonymous):

I think our formula is wrong...

OpenStudy (anonymous):

no way. Now you're just being silly.

OpenStudy (lgbasallote):

i tried lateral surface area too no cheese :/

OpenStudy (lgbasallote):

@Senior012 what is the unit of measurement for the choices? maybe that can explain the answers...

OpenStudy (saifoo.khan):

What's the point of fighting guys?

OpenStudy (anonymous):

Sorry the formula is : \[h = (SA - 2\pi r^2)/2\pi r\]The Numerator is : 1000−4(22)(3.5)(3.5)/7 = 1000 - 4(22)(0.5)(3.5) = 1000 - 77 = 923... And the Denominator is : 2(22)(3.5)/7=2(22)(0.5) = 22.. Now h = 923/22 = 41.95... Now the answer is pretty close to 41.97.. Isn't it....????

OpenStudy (lgbasallote):

oops oh yeah 2pi r^2

OpenStudy (lgbasallote):

i must have thought that the area of a circle is 2 pi r^2 for an odd reason

OpenStudy (anonymous):

Ha ha ha...Sometimes it happens...

OpenStudy (anonymous):

all hail the great @waterineyes for figuring it out---seriously.

OpenStudy (anonymous):

Discussion always solves the problem...